# MorphoSeg: An Uncertainty-Aware Deep Learning Method for Biomedical Segmentation of Complex Cellular Morphologies

### v1

@@ 5.5 Patch-Based Approach for High-Resolution Inference and Visualization
During a single image inference stage, the image is initially subdivided into patches of size $224\times 224$, similar to the data augmentation strategy presented in Section 4.2. This subdivision allows the model to focus on and accurately detect fine-grained variations at the boundaries of the Ntera-2 cells and their surrounding environment. Each patch undergoes a separate inference process. To preserve the global spatial context, we incorporate an overlapping strategy with a $56$-pixel margin. This overlap ensures continuity and minimizes edge artifacts. The inference results from the patches are then combined using an average pooling strategy and visualized as segmentation masks, shown in red in Figure 4. This end-to-end training and visualization strategy significantly facilitates the application of our framework, enabling researchers to use it without requiring further modifications.
6 Conclusion and Future Work
@@
The experimental results demonstrate that the proposed MorphoSeg outperforms existing benchmarks, including Cellpose and TransUNet. Our best-performing model, MorphoSeg with a hybrid encoder, showed a notable improvement in performance metrics, achieving a 6.23% increase in the Dice Similarity Coefficient (DSC). This statistic measures the similarity between the segmented cells and the ground truth. A reduction in HD95 (Hausdorff Distance at the 95th percentile) was also measured, indicating a decreased spatial discrepancy between the predicted and actual segmentation boundaries compared to the baseline configuration.
By providing comprehensive visualizations and complete code, we aim to contribute to not only advance further research in this emerging domain but also to serve as a practical solution for cell segmentation and visualization in diverse medical AI applications.
Future Work: Building on the proposed NT2 cells dataset, future research will address the challenge of segmenting complex cell morphologies, particularly those with substantial overlapping regions, which remain a significant obstacle to achieving accurate segmentation. Further efforts will focus on developing advanced deep-learning methods for biological AI to overcome these challenges.

### v2

@@ 5.6 Patch-Based Approach for High-Resolution Inference and Visualization
During a single image inference stage, the image is initially subdivided into patches of size $224\times 224$, similar to the data augmentation strategy presented in Section 4.2. This subdivision allows the model to focus on and accurately detect fine-grained variations at the boundaries of the Ntera-2 cells and their surrounding environment. Each patch undergoes a separate inference process. To preserve the global spatial context, we incorporate an overlapping strategy with a $56$-pixel margin. This overlap ensures continuity and minimizes edge artifacts. The inference results from the patches are then combined using an average pooling strategy and visualized as segmentation masks, shown in red in Figure 4 and Figure 5. This end-to-end training and visualization strategy significantly facilitates the application of our framework, enabling researchers to use it without requiring further modifications.
Figure 5: Zoomed-in visualization of Ntera-2 cell segmentation results on the testing dataset, including target cell annotations, predicted segmentation masks, and corresponding overlay outputs. As a detailed supplement to Figure 4, this comparison highlights the performance of MorphoSeg against TopoSeg and TransUNet. Notable differences between segmentation methods are observed, with some regions exhibiting incomplete segmentation, while others merge multiple cells into a single segmentation.
Figure 4 presents inference results with segmented masks overlaid on the original raw images. For comparison, we also show outputs from Cellpose, TransUNet, and TopoSeg using the same input images. It is evident that Cellpose fails to detect a significant number of cells, while TopoSeg produces unsatisfactory results with many spurious, thin lines. In contrast, MorphoSeg yields results comparable to TransUNet but with noticeable improvements in the fine details. Although the high resolution of the raw images can mask these subtle differences in the full-view images, the zoomed-in regions in Figure 5 clearly demonstrate the enhanced segmentation quality. These qualitative findings are consistent with the quantitative results reported in Table 1.
6 Conclusions and Future Works
@@
The experimental results demonstrate that the proposed MorphoSeg outperforms existing benchmarks, including Cellpose and TransUNet. Our best-performing model, MorphoSeg with a hybrid encoder, showed a notable improvement in performance metrics, achieving a 6.23% increase in the Dice Similarity Coefficient (DSC). This statistic measures the similarity between the segmented cells and the ground truth. A reduction in HD95 (Hausdorff Distance at the 95th percentile) was also measured, indicating a decreased spatial discrepancy between the predicted and actual segmentation boundaries compared to the baseline configuration. By providing comprehensive visualizations and complete code, we aim to contribute to not only advance further research in this emerging domain but also to serve as a practical solution for cell segmentation and visualization in diverse medical AI applications.
Limitations. The relatively small size of the annotated dataset (36 images), is a limitation and the data may not be sufficient for all research applications. We anticipate that incorporating a larger number of images could further improve model performance. Additionally, while our method is effective for the NT2 cell dataset, it may not scale seamlessly to all types of cell datasets; inappropriately handling systematic outliers could potentially reduce its effectiveness.
Future Work. Future work will focus on a broader range of deep learning architectures to further enhance performance and generalisability. This also includes applying sampling procedures within the latent space to further simplify training and potentially enhance the results. Quantification of the method performance with respect to uncertainties is another important aspect to study further.